# Probing a minimal dark gauge sector via microlensing of compact dark objects

### v1

@@ I Introduction
Within this point of view, we show that for $\mu\simeq 1$ eV, self-gravitating scalar field configurations have masses around $10^{-11}M_{\odot}$ and can play the role of massive compact halo objects (MACHOs). In such a scenario, the halo dynamics would be indistinguishable from that ina CDM halo, since N-body simulations of CDM are insensitive to particle mass granularity smaller than $10^{5}M_{\odot}$ Navarro et al. (2010). These scalar field mini-MACHOS could have formed in the early Universe via a mechanism analogous to the Affleck-Dine Mechanism known in particle physics Kasuya and Kawasaki (2000).
Static, spherically symmetric, self-gravitating systems minimally coupled to a scalar field with $U(1)$ gauge symmetry, known as charged boson stars, have been described in Jetzer and van der Bij (1989); Jetzer (1989); Pugliese et al. (2013); López and Alcubierre (2023). Recently, the dynamics of these self-gravitating configurations was extensively studied in Jaramillo et al. (2025). In this work, we reinterpret these configurations as dark matter structures within the DSM framework. The resulting solutions are everywhere regular and horizonless, representing solitonic equilibria of the model. In contrast, black hole configurations with nontrivial scalar field profiles exist only for self-interacting scalars under specific resonance conditions Herdeiro and Radu (2020); Hong et al. (2020). Such solutions can be continuously connected to the self-interacting (Q-ball–like) boson star sequence when a small event horizon is introduced. However, full 3+1 numerical evolutions show that they develop a non-axisymmetric instability Nicoules et al. (2025), while their horizonless counterparts remain stable Jaramillo et al. (2025). Therefore, within this theory, the viable static gravitational solitons that could serve as MACHO dark matter candidates are the Reissner–Nordström black holes and the gauged boson stars. Throughout the paper we adopt geometric units ($c=1=G$) except where stated otherwise. In this units, the coupling parameter $q$ then has units of inverse length and can be normalized by the scalar field mass $\mu$, so that ratio $q/\mu$ becomes dimensionless.
@@ II Gauge dark matter solitons
Under these assumptions for the spacetime and the spin-0 and spin-1 fields, the system reduces to four coupled ordinary differential equations for the functions $V(r)$, $\phi(r)$, $F_{0}(r)$ and $F_{1}(r)$, see e.g. Appendix B in Jaramillo et al. (2025). We solve the system numerically, noting first that the mass of the scalar field $\mu$ can be factored out as the characteristic length scale of the system, and all dimensional quantities rescale with appropriate powers $\mu$. We solve the system using a spectral method with a compactified radial coordinate $r$ and a Newton-Raphson algorithm, as described in Alcubierre et al. (2022), and employ 24 spectral coefficients for all solutions presented.
@@
Figure 2: Summary of current microlensing limits on MACHO dark matter fraction as a function of lens mass. The solid black lines mark the $95\%$ confidence level exclusion boundaries from recent surveys: Niikura et al. Niikura et al. (2019) covers the sub-lunar window around $10^{-11}-10^{-6}\,M_{\odot}$, Smyth et al. Smyth et al. (2020) covers the Earth–Jupiter range $10^{-6}-10^{-3}\,M_{\odot}$, Wyrzykowski et al. Wyrzykowski et al. (2011) covers the stellar range $\sim 10^{-1}-10\,M_{\odot}$, and Blaineau et al. Blaineau et al. (2022) covers the intermediate range $10-10^{3}\,M_{\odot}$. The red-shaded regions within these curves are excluded.
@@ III Constraints on the free $\mu$ and $q$ from microlensing limits
Current microlensing observations have imposed tight constraints on the fraction of dark matter that may be due to MACHOs. As summarized in Fig. 2, these limits depend on the mass of the compact objects $\mathcal{M}$. Microlensing observations have effectively exclude MACHOs with $\mathcal{M}>10^{-11}M_{\odot}$ as the dominant dark matter. In the sub-lunar to Earth-mass range ($10^{-11}\lesssim\mathcal{M}/M_{\odot}\lesssim 10^{-6}$), the Subaru Hyper Suprime-Cam (HSC) survey of M31 by Niikura et al. Niikura et al. (2019) detected only one candidate event, placing strong upper limits on the MACHO fraction of dark matter and effectively closing the so-called “lunar-mass window”. At lower masses ($10^{-16}\lesssim\mathcal{M}/M_{\odot}\lesssim 10^{-10}$), Smyth et al. Smyth et al. (2020) reanalyzed the same HSC data, incorporating finite-source-size effects. They found that these effects significantly weaken constraints at the smallest masses, suggesting that asteroid-mass MACHOs could still constitute all of the dark matter. At solar masses and above, OGLE observations toward the Magellanic Clouds, particularly those by Wyrzykowski et al. Wyrzykowski et al. (2011), show that the detected microlensing events are consistent with known stellar populations, thereby ruling out MACHOs with $\mathcal{M}\sim 0.1-10\,M_{\odot}$ as more than a small fraction of the dark halo. For heavier objects in the intermediate range ($\mathcal{M}\sim 10-10^{3}\,M_{\odot}$), Blaineau et al. Blaineau et al. (2022) combined MACHO and EROS datasets, finding that such black hole–like MACHOs can account for at most about $15\%-50\%$ of the dark matter, depending on mass. Overall, microlensing surveys have largely ruled out MACHOs as the dominant form of dark matter, leaving room primarily for mini-MACHOs in the asteroid-mass regime.

### v2

@@ I Introduction
Within this point of view, we show that for $\mu\simeq 1$ eV, self-gravitating scalar field configurations have masses around $10^{-11}M_{\odot}$ and can play the role of massive compact halo objects (MACHOs). In such a scenario, the halo dynamics would be indistinguishable from that in a CDM halo, since N-body simulations of CDM are insensitive to particle mass granularity smaller than $10^{5}M_{\odot}$ Navarro et al. (2010). These scalar field mini-MACHOS could have formed in the early Universe via a mechanism analogous to the Affleck-Dine Mechanism known in particle physics Kasuya and Kawasaki (2000). Dark neutron stars have recently been investigated within a related framework Litterer and Rosa (2025), while Ref. Boos and Hu (2026) analyzes gravitational microlensing by a class of static, spherically symmetric nonsingular black holes. Related Higgs-portal constructions with a spontaneously broken global U(1) symmetry have also been explored Ng et al. (2014). More recently, it has been shown that nonminimal couplings to gravity may produce demagnification signatures in microlensing light curves Zhang (2025a).
Static, spherically symmetric, self-gravitating systems minimally coupled to a scalar field with $U(1)$ gauge symmetry, known as charged boson stars, have been described in Jetzer and van der Bij (1989); Jetzer (1989); Pugliese et al. (2013); López and Alcubierre (2023). Recently, the dynamics of these self-gravitating configurations was studied in Jaramillo et al. (2025). In this work, we reinterpret these configurations as dark matter structures within the DSM framework. The resulting solutions are everywhere regular and horizonless, representing solitonic equilibria of the model. In contrast, black hole configurations with nontrivial scalar field profiles exist only for self-interacting scalars under specific resonance conditions Herdeiro and Radu (2020); Hong et al. (2020). Such solutions can be continuously connected to the self-interacting (Q-ball–like) boson star sequence when a small event horizon is introduced. However, full 3+1 numerical evolutions show that they develop a non-axisymmetric instability Nicoules et al. (2025), while their horizonless counterparts remain stable Jaramillo et al. (2025). Therefore, within this theory, the viable static gravitational solitons that could serve as MACHO dark matter candidates are the Reissner–Nordström black holes and the gauged boson stars. Throughout the paper we adopt geometric units ($c=1=G$) except where stated otherwise. In these units, the coupling parameter $q$ then has units of inverse length and can be normalized by the scalar field mass $\mu$, so that ratio $q/\mu$ becomes dimensionless.
@@ II Gauge dark matter solitons
Under these assumptions for the spacetime and the spin-0 and spin-1 fields, the system reduces to four coupled ordinary differential equations for the functions $V(r)$, $\phi(r)$, $F_{0}(r)$ and $F_{1}(r)$, see e.g. Appendix B in Jaramillo et al. (2025) (or Zhang (2025b) for a more general treatment in a non-relativistic limit). We solve the system numerically, noting first that the mass of the scalar field $\mu$ can be factored out as the characteristic length scale of the system, and all dimensional quantities rescale with appropriate powers $\mu$. We solve the system using a spectral method with a compactified radial coordinate $r$ and a Newton-Raphson algorithm, as described in Alcubierre et al. (2022), and employ 24 spectral coefficients for all solutions presented.
@@
Figure 2: Summary of current microlensing limits on the MACHO dark matter fraction as a function of lens mass. The solid black line marks the combined 95% confidence level exclusion boundary obtained from recent surveys. Niikura et al. Niikura et al. (2019) constrain the sub-lunar window around $10^{-11}-10^{-6}\,M_{\odot}$, Smyth et al.Smyth et al. (2020) probe the Earth–Jupiter range $10^{-6}-10^{-3}\,M_{\odot}$, Wyrzykowski et al. Wyrzykowski et al. (2011) constrain the stellar regime $\sim 10^{-1}-10\,M_{\odot}$, and Blaineau et al. Blaineau et al. (2022) cover the intermediate range $10-10^{3}\,M_{\odot}$. Updated OGLE analyses combining the results of Refs. Mróz et al. (2024a, b) further impose global limits, excluding MACHOs contributing more than 1% of the dark matter in the mass range $10^{-8}-10\,M_{\odot}$, and more than 10% in the broader interval $10^{-9}-10^{3}\,M_{\odot}$. The red-shaded regions above the exclusion boundary are therefore ruled out.
@@ III Constraints on the free $\mu$ and $q$ from microlensing limits
Current microlensing observations have imposed tight constraints on the fraction of dark matter that may be due to MACHOs. As summarized in Fig. 2, these limits depend on the mass of the compact objects $\mathcal{M}$. Microlensing observations have effectively excluded MACHOs with $\mathcal{M}>10^{-11}M_{\odot}$ as the dominant dark matter. In the sub-lunar to Earth-mass range ($10^{-11}\lesssim\mathcal{M}/M_{\odot}\lesssim 10^{-6}$), the Subaru Hyper Suprime-Cam (HSC) survey of M31 by Niikura et al. Niikura et al. (2019) detected only one candidate event, placing strong upper limits on the MACHO fraction of dark matter and effectively closing the so-called “lunar-mass window”. At lower masses ($10^{-16}\lesssim\mathcal{M}/M_{\odot}\lesssim 10^{-10}$), Smyth et al. Smyth et al. (2020) reanalyzed the same HSC data, incorporating finite-source-size effects. They found that these effects significantly weaken constraints at the smallest masses, suggesting that asteroid-mass MACHOs could still constitute all of the dark matter. At solar masses and above, OGLE observations toward the Magellanic Clouds, particularly those by Wyrzykowski et al. Wyrzykowski et al. (2011), show that the detected microlensing events are consistent with known stellar populations, thereby ruling out MACHOs with $\mathcal{M}\sim 0.1-10\,M_{\odot}$ as more than a small fraction of the dark halo. For heavier objects in the intermediate range ($\mathcal{M}\sim 10-10^{3}\,M_{\odot}$), Blaineau et al. Blaineau et al. (2022) combined MACHO and EROS datasets, finding that such black hole–like MACHOs can account for at most about $15\%-50\%$ of the dark matter, depending on mass. Overall, microlensing surveys have largely ruled out MACHOs as the dominant form of dark matter, leaving room primarily for mini-MACHOs in the asteroid-mass regime.